OpenStudy (anonymous):

The sum of two numbers is -33. One number is 3 more than twice the other. Find the larger of the two numbers.

OpenStudy (mertsj):

x= first number -33-x=second number x=2(-33-x)+3

OpenStudy (mertsj):

Solve tat and you'll have it.

OpenStudy (anonymous):

would -12 be the answer?

OpenStudy (anonymous):

Let x = first number Let y = second number x+y=-33 x=2y+3 So the larger number is 'x' because it's more than twice the value of 'y'.

OpenStudy (mertsj):

yes. -12 would be the answer.

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

(2y+3)+y=-33 3y+3=-33 -3 -3 3y=-36 /3 /3 y=-12 Yep! :)

OpenStudy (anonymous):

What about this problem

OpenStudy (anonymous):

Find two numbers whose sum is 42 and whose difference is 16.

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

x+y=42 x-y=16

OpenStudy (mertsj):

Add the two equations to eliminate y and solve for x

OpenStudy (anonymous):

would it be 13 and 23?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

If you add those two equations together, you get: 2x=58 /2 /2 x=29